Minutes

Meeting Date File
September 22, 2017

216. BAs flooding issues and their mitigation, REX Isolde RF improvement, results and consolidation plans, Report on LINAC3 amplifier failure, Mandate for AD and ISOLDE technical coordinators

September 15, 2017

215. aC coating in the YETS, Isolde: plans and status for Module 4, Isolde radiation issues with 94Rb beams

September 1, 2017

214. Cooling issues in the NA, Consolidation of the AD ring quadrupoles, Report from EACM on LS2 activities, Crab Cavity Operational Scenario, Update on Gbar readiness for beam permit

August 18, 2017

213. Update on the injector TS, 2017 schedule update: possible extension of 2 weeks in ISOLDE, n_TOF, EA

August 11, 2017

212. ISOLDE High Voltage power converter consolidation plans, Report on the recent SPS dump kicker trigger issues and mitigation plans, Update on the injector TS

August 4, 2017

211. INTC report n_TOF, INTC report ISOLDE, Update on the availability of the MORPURGO magnet, Injector technical stop

July 21, 2017

210. SPS access system upgrade in LS2. Report of International Advisory Panel of the HIE-ISOLDE project. TIDVG#3 post mortem leak detection. Access system in b. 361.

July 14, 2017

209. EYETS coordination post-mortem. Summary of the International Review on PS Powering Strategy. Plans for tests on SPS crystal extraction and on other potential extraction loss reduction measures. Consolidation of the Linac3 HLRF system. SPS ZS tank 2 alignment issues: solution to the losses.

July 7, 2017

208. Non-LHC beam performances afer Linac4 connection, Status and future prospects of distributed optical fibre radiation and temperature sensing at CERN, Update on the master schedule of YETS 2017/18, Reliability run Linac4

June 23, 2017

207. North area cooling and other specific topics, Report from the SPSC, New functional specification for impedance Shielding of new SPS Sector Valves

June 16, 2017

206. Follow-up of a selection of IEFC actions, SPS-TIDVG status report, Status of building 359 false floor

June 2, 2017

205. Plans for COLDEX in 2017, Report on the false floor incident in Building 359, Lessons learned from PSB repowering, Status of the PSB Ring 2 C02 cavity issue

May 19, 2017

204. Plans for MEDICIS in 2017, Outcome of the injector MD days, Report on East Area Renovation project, SPS TIDVG update, Consolidation ot the SPS intercom

May 12, 2017

203. Summary of the PSB & PS RF by-pass review, Outcome of the AWWAKE review, Crab cavities: outcome of the review, status and commissioning plans, Mandate for the North Area consolidation study

April 28, 2017

202. Report from the SPSC, Vacuum specification for components to be installed in the accelerators, Plans for HiRadMat in 2017, Update on Linac2 vaccum leak, Status of HIE-ISOLDE

April 21, 2017

201. PS cable identification status, Outcome of the GBAR safety review

April 7, 2017

200. Update and plans for CLEAR, YETS planning update 2017-2018, Maintenance of Linac2 RF during the EYETS 2017, Activities on the Linac4 LLRF for the reliability run in 2017, Status of Linac2 vaccuum

March 24, 2017

199. Plans for the Xenon ions throughout the injector complex in 2017, status of the new POPS capacitor container, impact on the PSB restart of delays in equipment repowering.

March 17, 2017

198. PS 40/80 MHz spare situation and needs, LINAC4 Reliability Run readiness and plans, Status of the SPS tunnel transfer lines (TT0, TT20), PSB repowering status update, Matrix changes in the NA

March 10, 2017

197. BioLEIR status and plans, Status of the POPS containers and POPS recommissioning plans, Report on Linac2 proton source tests, Report on the 66kV power cut, Issue with SPS SEM-grid upstream of the TIDVG

Pages

You are here